OpenStudy (abbycross167):

Can someone please check the answer to my question? Please? Choose an equation in standard form of the line that passes through the point (-1, -4) and has a slope of 2. a. -2x + y = -2 b. -3y = x – 9 c. –x – y = 9 d. x – 3y = -9

OpenStudy (abbycross167):

@kittiwitti1

OpenStudy (kittiwitti1):

I was about to sign off :|

OpenStudy (abbycross167):

Oh, I am very sorry, Don't worry about it! I'll figure it out

OpenStudy (kittiwitti1):

Too late. ○ D ○ point slope:\[y-y_{1}=m(x-x_{1})\]putting the coordinates in,\[y-(-4)=2(x-(-1))\]\[y+4=m(x+1)\rightarrow y+4=2x+2\rightarrow y=2x-2\]

OpenStudy (kittiwitti1):

Standard form: -2x+y=-2

OpenStudy (kittiwitti1):

Now I've got to go~ • v •)/

OpenStudy (abbycross167):

okay, thank you very much!!!